OpenStudy (anonymous):

If Ari, Ben, and Carol each choose a number at random between 1 and 5, inclusive, what is the probabilty that all of them choose different numbers? (a) 1/125 (b) 1/25 (c) 1/5 (d) 3/5 (e) 12/25

OpenStudy (anonymous):

The probability is\[ \frac{_5P_3}{5^3}=\frac{5\cdot4\cdot3}{5^3} \]I'm sure you can work from there

OpenStudy (anonymous):

thanks